OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(-2, 7) B(-2,-2)? A) y=-2 B) Y= x C) 3x=5 D) -2x=7y

OpenStudy (anonymous):

Because your 2 points creates a vertical line, your line that is parallel to it must have only the variable x in the equation.

OpenStudy (anonymous):

Okay, so how would i do this

OpenStudy (solomonzelman):

the slope will be the same, the y intercept will only change.

OpenStudy (solomonzelman):

what is the slope of A(-2, 7) B(-2,-2)?

OpenStudy (anonymous):

I have no idea... I'm sorry i really bad at this

OpenStudy (solomonzelman):

\[\frac{y_1-y_2}{x_1-x_2}=slope\] plug in your values.

OpenStudy (anonymous):

okay so Y1-y2= -2-(-2) and X1-X2 =7-(-2) is this right??

OpenStudy (solomonzelman):

Yes!

OpenStudy (anonymous):

Okay so what do i do now?

OpenStudy (solomonzelman):

so you see that \[\frac{0}{9}=slope--->slope=0\]

OpenStudy (solomonzelman):

find what y equals to, the y axis that this (horizontal) line is on.

OpenStudy (anonymous):

I hope you realize your what you did was flipped and that your slope is undefined

OpenStudy (solomonzelman):

@wastetime, you are right, SORRY!

OpenStudy (anonymous):

Im so confused

OpenStudy (solomonzelman):

it is 9/0

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

@wastetime, take over please.....

OpenStudy (anonymous):

When your slope is undefined or in other words some number over 0, your line is only a vertical line (straight up and down). The equation of a vertical line is x = some number so it corresponds to all y values but only has that one x-value. From my first post, the only variable in your answer should be x.

OpenStudy (anonymous):

okay so it would be (B) Y= x

OpenStudy (anonymous):

Y is another variable. So (B) Y=x is incorrect.

OpenStudy (anonymous):

oh my.. Okay

OpenStudy (anonymous):

3x=5?

OpenStudy (anonymous):

Yes it is!

OpenStudy (anonymous):

I get it now! I misread it (: would you mind helping me with more?

OpenStudy (anonymous):

Sure. You can post them or private message them if you like